Calculate the boiling point of a solution made from 227 g of MgCl2 dissolved in 700. g of water. What is the boiling point of the solution? Kb = 0.512° C/m.

1
Expert's answer
2022-04-15T04:10:03-0400

ΔTb = iKbCm

i(MgCl2) = 3


Cm = nsolute / msolvent

Cm = n(MgCl2) / m(H2O) = (m(MgCl2))/Mr(MgCl2))) / m(H2O) = (227 g / 94 g/mol) / 0.700 kg = 3.45 mol/kg


ΔTb = 3 * 0.512° C/m * 3.45 mol/kg = 5.298°C

Tb (solution) = ΔTTb (pure solvent)

Tb = ΔT + Tb(H2O) = 5.298°C + 100°C = 105.298°C